# Exceptional-point-controlled mode interaction in three-dimensional microcavities represented by generalized Husimi functions

### v1

@@ VI Acknowledgments
The authors would like to thank the Scientific Computing and Data Analysis Section at OIST. SNC acknowledges support from the TU Chemnitz Visiting Scholar Program. We thank Takahisa Harayama, Roland Ketzmerick and Arnd Bäcker for discussions. This work was partly supported by OIST Graduate University.
Supplementary Material
@@ Appendix A Husimi functions for TE modes
For $\eta=0$ (see Fig. 10a) only a single intensity maximum is visible in the Husimi function. For $\eta>0$ the single maximum splits in two maxima, and the angle to $\sin\chi_{xy}$ increases with $\eta$, which is shown in Figs. 10b-c.
Note that the different boundary conditions for TE polarization results in a mode morphology which is slightly different to that of TM modes (compare Fig. 4). This leads to a slight washing out of the Husimi function of $\Omega^{\mathrm{TE}}_{3}$ mode in Fig. 10b, which is not present for its TM mode counterpart in Fig. 4b.
[…]
Appendix B Cone Unwrapping
@@
$\varphi^{\prime}&=\varphi\sin\alpha\quad\text{and}\\ z^{\prime}&=z/\cos\alpha,\quad\end{aligned}$ (18)
with an angle scale factor $\sin\alpha$ and a height scale factor $1/\cos\alpha$. The new coordinates $(\varphi^{\prime},z^{\prime})$ can be understood as a 2D polar representation of points on the cylinder surface. Cartesian coordinates $(s,\tilde{z})$ on the flat cone mantle can then be easily obtained using the usual transformation
$s&=z^{\prime}\cos\varphi^{\prime}\quad\text{and}\\ \tilde{z}&=z^{\prime}\sin\varphi^{\prime}.\quad\end{aligned}$ (19)
An example is shown in Fig. 11, where the transformation is applied to points on the surface of cone with an opening angle of $\alpha=0.04\pi$. The yellow (purple) line corresponds to the boundary of the top (bottom) surface of a slanted cylinder of height $h=4$ and a deformation $\delta=0.5$. Figure 11a shows points on the surface for various $z$ values, mapped to a 2D plane in Fig. 11b.
Using this method, the rotationally symmetric wave functions (such as in Fig. III.4, upper panels) can be mapped to a flat, 2D plane. A phasor is applied to the numerically calculated 2D field function $\psi_{m}(r,z)$ with angular momentum $m$ such that
@@
The aforementioned transformation is then applied to the field function on the boundary. The resulting field on the surface for an example of $\Omega_{2}(\delta=0.1)$ in Fig. III.4e is shown in Fig. 12a.
Appendix C Spatial Dependence of Husimi Functions
Figure 12: (a) Electric field of the $\Omega_{2}$ mode with $(\rho,m,\eta)=(1,6,2)$ of Fig. III.4 on the cylinder surface, the black dashed line notes the taken coherent state positions for Fig. C. (b-d) Husimi functions depending on coherent state position. The positions are $s=0$ and (b, dot) $z/R_{0}=1$, (c, diamond) $z/R_{0}=2.5$, and (d, cross) $z/R_{0}=4$. The black dashed line denotes the expected Husimi function maximum position from Eq. (21).
When cavity deformation is increased, the modes lose the mirror symmetry with respect to the $xy$-plane. The qualitative mode morphology now depends on the location in the cavity which also implies a spatial dependence of the Husimi function. This is examined in Fig. 12. Figure 12a shows the unrolled surface of the $(\rho,\eta)=(1,2)$ mode. Three different coherent state positions are selected (see dot, diamond, and cross markers) and the Husimi function is calculated. The resulting phase spaces are shown in Fig. 12b-d. The Husimi function calculated at the bottom of the cavity (see black dot) shows the signature of a mode with nonzero $\eta$, with two distinct Husimi maxima, see Fig. 12b. The phase space in the cavity center (diamond) mainly measures a single broad intensity maximum, which results in a single intensity maximum at $\sin\chi_{z}\approx 0$, see Fig. 12c. This is a result of the finite size of the coherent state. The top coherent state (cross) is located in an area where the field intensity is very low. The field function is dominated by the weak remains of the top intensity maximum, leading to a feint single maximum signature in the Husimi function, see Fig. 12d.
To gain deeper insight into the spatial dependence of the Husimi function we now vary the coherent state position continuously. We select two different paths on the cavity surface: (i, horizontal dashed line in Fig. 12a) $\tilde{z}=2.5$ with $\varphi\in[0,2\pi/m]$, covering a period in $\mathrm{e}^{im\varphi}$ (see Eq. (3)); and (ii, vertical dashed line in Fig. 12a) $\varphi=0$ with $\tilde{z}\in[0,h/\cos\alpha]$. We only consider the Husimi function values for a fixed value of
@@ Appendix C Spatial Dependence of Husimi Functions
{tblr}
c—[dashed]c—[dashed]c $\delta=0$ $\delta=0.1$
[dashed] $\eta=2$ Refer to caption Refer to caption
Figure 13: Husimi functions along path (i) (see Fig. 12a) and $\sin\chi_{xy}$ according to Eq. (21) for modes with $\eta=2$. The deformations are (a) $\delta=0$ and (b) $\delta=0.1$.
The results for path (i) shown in Fig. C. The Husimi function for $\delta=0$ (see Fig. III.4e) in Fig. Ca does not depend on the azimuthal coordinate on the surface, as is expected for a system with rotational symmetry. The same is true for the deformed case in Fig. Cb (compare Fig. III.4f).
Path (ii) yields more interesting results, which are shown in Fig. C. The morphology of the mode with $\eta=0$ showed a significant change in its morphology when the deformation is increased in Fig. III.4c,d. This is also visible in phase space, where the intensitiy maximum shifts from $z\approx h/2$ in Fig. Ca to $z\approx h/5$ in Fig. Cb. For $\eta=2$ in Figs. Cc,d the change in the Husimi function is even more pronounced. In the unperturbed case, the field intensity maxima positions in real space (compare Fig. III.4e) can be identified by the nonzero Husimi function intensities $\sin\chi_{z}=0$. For increased deformation the field intensity maximum at $z\approx h/2$ (cf. Fig. III.4f) is clearly visible in the Husimi function as a broad intensity maximum. At lower height the field is squeezed in $z$ direction, which is visible as a bimodal distribution in the Husimi function, as the real-space checkerboard wave function mimicks a mode with higher $\eta$. For $z>h/2$ the Husimi function signature disappears to to the low field intensity. When the excitiation in $z$ is high, e.g. $\eta=7$, the Husimi function intensity is confined to $\sin\chi_{z}=\pm 0.4$ in Fig. Ce. For the deformed cylinder in Fig. Cf the distance in $\sin\chi_{z}$ of two intensity maxima decreases with $z$. This is the result of a similar squeeze of the wave function as in Fig. Cd, but here the maximum at the top is significantly smaller in real space.
Interestingly, the orthogonality constraints on the modes discussed in Sec. III.4 are also visible in the Husimi function signatures. In the deformed cavity, the modes with higher $\eta$ occupy different regions in phase space. The intensity Fig. Cd has a distinct minimum at $(z,\sin\chi_{z})\approx(2,0)$, which is precisely where the maximum in Fig. Cb is located. The Husimi function for $\eta=7$ in Fig. Cf appears to encircle the aforementoned phase space structures.
{tblr}
c—[dashed]c—[dashed]c $\delta=0$ $\delta=0.1$
[…]
Figure 14: Husimi functions along path (ii) (cf. Fig. 12a) and $\sin\chi_{xy}$ according to Eq. (21) for modes with $\rho=1$ and various $\eta$ (cf. Fig. III.4). Where the values are taken in real and phase space is shown in Fig. 12.
@@ Appendix D Additional ARCs
{tblr}
c—[dashed]c—[dashed]c—[dashed]c—[dashed]c $\delta=0$ $\delta=0.075$ $\delta=0.12$ $\delta=0.16$
[dashed] upper branch Refer to caption Refer to caption Refer to caption Refer to caption
[dashed] center branch Refer to caption Refer to caption Refer to caption Refer to caption
[dashed] lower branch Refer to caption Refer to caption Refer to caption Refer to caption
[…]
There are many types of ARCs evident in Fig. 6. The TM-TE type ARC shown in Fig. 7 is one possible combination of mode polarizations. Crossings involving more than one mode also exist, one if which is shown in Fig. D. It involves two TE and one TM polarized modes. Noticeably, only the upper and the lower branch exhibit mode hybridization at the ARC at $\delta=0.12$; this is also visible in the crossing of imaginary components of $\Omega$ in Fig. Db. The middle branch experiences an enhanced lifetime, which is caused by the mode pattern resembling a mode with higher $\eta$ than the unperturbed mode. Modes with high $\eta$ generally have longer lifetimes. The behavior of an ARC involving $n$ modes is described by an $n\times n$ Hamiltonian [39, 70].
@@ Appendix E Husimi functions at the EP
[…]
Figure 16: Top row: Electric (red) and magnetic (blue) fields. Bottom row: Associated average Husimi functions close to the EP at ${\delta=0.222}$ and ${n=2.435}$.

### v2

@@ V Conclusion
Experimentally, our findings could be accessed via a host of platforms, as ARCs between TM and TE-polarized modes are very common in 3D cavities. For any given refractive index and wavelength, a suitable ARC can be chosen to which the system geometry can be tailored. The only crucial aspect is, that the contributing resonances should have a high $\eta$, hence the system needs to have dimensions $h/R_{0}>1$. The presence of the ARC (or the EP) could be confirmed by measuring Stokes’ parameters, as the resonances are only unpolarized in the vicinity of the ARC. We note that the formalism can also be applied to cavities with nonlinearities, such as microlasers with gain [12, 78].
@@ VI Acknowledgments
T.S.R. would like to thank the Scientific Computing and Data Analysis Section at OIST for access and support. S.N.C. acknowledges support from the TU Chemnitz Visiting Scholar Program. We thank Takahisa Harayama, Jung-Wan Ryu, Chang-Hwan Yi, Roland Ketzmerick, and Arnd Bäcker for discussions, M.H. in particular for the fruitful exchange at Waseda University and at the Center for Theoretical Physics of Complex Systems, Daejeon. This work was partly supported by OIST Graduate University. S.L. acknowledges support from the Japan Society for the Promotion of Science (JSPS) KAKENHI through Grant-in-Aid for Scientific Research (C) Grant No. 23K04617.
@@ Appendix A Husimi functions for TE modes
For $\eta=0$ (see Fig. 11(a)) only a single intensity maximum is visible in the Husimi function. For $\eta>0$ the single maximum splits in two maxima, and the angle to $\sin\chi_{xy}$ increases with $\eta$, which is shown in Figs. 11(b-c).
Note that the different boundary conditions for TE polarization results in a mode morphology which is slightly different to that of TM modes (compare Fig. 5). This leads to a slight washing out of the Husimi function of $\Omega^{\mathrm{TE}}_{3}$ mode in Fig. 11(b), which is not present for its TM mode counterpart in Fig. 5(b).
Figure 11: Field intensity (left) and Husimi functions (right) for TE modes, with $(\rho,m)=(0,6)$ and various $\eta$. The frequencies are (a) $(\eta,\Omega)=(0,3.8377-i\,0.0070)$, (b) $(\eta,\Omega)=(3,3.9945-i\,0.0015)$ and (c) $(\eta,\Omega)=(5,4.1901-i\,0.0006)$ The phase space functions are shown for $\sin\chi_{xy}>0$, the rest is given by symmetry.
Appendix B Unwrapping the cone
@@
$\begin{aligned} \varphi^{\prime}&=\varphi\sin\alpha\quad\text{and}\\ z^{\prime}&=z/\cos\alpha,\quad\end{aligned}$ (18)
[…]
$\begin{aligned} s&=z^{\prime}\cos\varphi^{\prime}\quad\text{and}\\ \tilde{z}&=z^{\prime}\sin\varphi^{\prime}.\quad\end{aligned}$ (19)
An example is shown in Fig. 12, where the transformation is applied to points on the surface of cone with an opening angle of $\alpha=0.04\pi$. The yellow (purple) line corresponds to the boundary of the top (bottom) surface of a slanted cylinder of height $h=4$ and a deformation $\delta=0.5$. Fig. 12(a) shows points on the surface for various $z$ values, mapped to a 2D plane in Fig. 12(b).
Using this method, the rotationally symmetric wave functions (such as in Fig. 6, upper panels) can be mapped to a flat, 2D plane. A phasor is applied to the numerically calculated 2D field function $\psi_{m}(r,z)$ with angular momentum $m$ such that
@@
The aforementioned transformation is then applied to the field function on the boundary. The resulting field on the surface for an example of $\Omega_{2}(\delta=0.1)$ in Fig. 6(e) is shown in Fig. 13(a).
Appendix C Spatial dependence of the Husimi function
Figure 13: (a) Electric field of the $\Omega_{2}$ mode with $(\rho,m,\eta)=(1,6,2)$ of Fig. 6 on the cylinder surface, the black dashed line notes the taken coherent state positions for Fig. 15. (b-d) Husimi functions depending on coherent state position. The positions are $s=0$ and (b, dot) $z/R_{0}=1$, (c, diamond) $z/R_{0}=2.5$, and (d, cross) $z/R_{0}=4$. The black dashed line denotes the expected Husimi function maximum position from Eq. (21).
When cavity deformation is increased, the modes lose the mirror symmetry with respect to the $xy$-plane. The qualitative mode morphology now depends on the location in the cavity which also implies a spatial dependence of the Husimi function. This is examined in Fig. 13. Figure 13a shows the unrolled surface of the $(\rho,\eta)=(1,2)$ mode. Three different coherent state positions are selected (see dot, diamond, and cross markers) and the Husimi function is calculated. The resulting phase spaces are shown in Fig. 13(b-d). The Husimi function calculated at the bottom of the cavity (see black dot) shows the signature of a mode with nonzero $\eta$, with two distinct Husimi maxima, see Fig. 13(b). The phase space in the cavity center (diamond) mainly measures a single broad intensity maximum, which results in a single intensity maximum at $\sin\chi_{z}\approx 0$, see Fig. 13(c). This is a result of the finite size of the coherent state. The top coherent state (cross) is located in an area where the field intensity is very low. The field function is dominated by the weak remains of the top intensity maximum, leading to a feint single maximum signature in the Husimi function, see Fig. 13(d).
To gain deeper insight into the spatial dependence of the Husimi function we now vary the coherent state position continuously. We select two different paths on the cavity surface: (i, horizontal dashed line in Fig. 13(a)) $\tilde{z}=2.5$ with $\varphi\in[0,2\pi/m]$, covering a period in $\mathrm{e}^{im\varphi}$ (see Eq. (3)); and (ii, vertical dashed line in Fig. 13(a)) $\varphi=0$ with $\tilde{z}\in[0,h/\cos\alpha]$. We only consider the Husimi function values for a fixed value of
@@ Appendix C Spatial dependence of the Husimi function
$\delta=0$ $\delta=0.1$
[dashed]
$\eta=2$
Refer to caption Refer to caption
Figure 14: Husimi functions along path (i) (see Fig. 13(a)) and $\sin\chi_{xy}$ according to Eq. (21) for modes with $\eta=2$. The deformations are (a) $\delta=0$ and (b) $\delta=0.1$.
$\delta=0$ $\delta=0.1$
[dashed]
$\eta=0$
Refer to caption Ω= 4.820 - i  0.038 Refer to caption Ω= 4.487 - i  0.032
[dashed]
$\eta=2$
Refer to caption Ω= 4.854 - i  0.035 Refer to caption Ω= 4.630 - i  0.028
[dashed]
$\eta=7$
Refer to caption Ω= 5.185 - i  0.013 Refer to caption Ω= 5.002 - i  0.015
Figure 15: Husimi functions along path (ii) (cf. Fig. 13(a)) and $\sin\chi_{xy}$ according to Eq. (21) for modes with $\rho=1$ and various $\eta$ (cf. Fig. 6). Where the values are taken in real and phase space is shown in Fig. 13.
The results for path (i) are shown in Fig. 14. The Husimi function for $\delta=0$ (see Fig. 6(e)) in Fig. 14(a) does not depend on the azimuthal coordinate on the surface, as is expected for a system with rotational symmetry. The same is true for the deformed case in Fig. 14(b) (compare Fig. 6(f)).
Path (ii) yields more interesting results, which are shown in Fig. 15. The morphology of the mode with $\eta=0$ showed a significant change in its morphology when the deformation is increased in Figs. 6(c,d). This is also visible in phase space, where the intensitiy maximum shifts from $z\approx h/2$ in Fig. 15(a) to $z\approx h/5$ in Fig. 15(b). For $\eta=2$ in Figs. 15(c,d) the change in the Husimi function is even more pronounced. In the unperturbed case, the field intensity maxima positions in real space (compare Fig. 6(e)) can be identified by the nonzero Husimi function intensities $\sin\chi_{z}=0$. For increased deformation the field intensity maximum at $z\approx h/2$ (cf. Fig. 6(f)) is clearly visible in the Husimi function as a broad intensity maximum. At lower height the field is squeezed in $z$ direction, which is visible as a bimodal distribution in the Husimi function, as the real-space checkerboard wave function mimicks a mode with higher $\eta$. For $z>h/2$ the Husimi function signature disappears to to the low field intensity. When the excitiation in $z$ is high, e.g. $\eta=7$, the Husimi function intensity is confined to $\sin\chi_{z}=\pm 0.4$ in Fig. 15(e). For the deformed cylinder in Fig. 15(f) the distance in $\sin\chi_{z}$ of two intensity maxima decreases with $z$. This is the result of a similar squeeze of the wave function as in Fig. 15(d), but here the maximum at the top is significantly smaller in real space.
Interestingly, the orthogonality constraints on the modes discussed in Sec. III.4 are also visible in the Husimi function signatures. In the deformed cavity, the modes with higher $\eta$ occupy different regions in phase space. The intensity Fig. 15(d) has a distinct minimum at $(z,\sin\chi_{z})\approx(2,0)$, which is precisely where the maximum in Fig. 15(b) is located. The Husimi function for $\eta=7$ in Fig. 15(f) appears to encircle the aforementoned phase space structures.
@@ Appendix D Additional ARCs
$\delta=0$ $\delta=0.075$ $\delta=0.12$ $\delta=0.16$
[dashed]
upper branch
Refer to caption Refer to caption Refer to caption Refer to caption
[dashed]
center branch
Refer to caption Refer to caption Refer to caption Refer to caption
[dashed]
lower branch
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: (a,b) ARC involving two TE modes and one TM mode and (c-n) wave functions at characteristic points of the ARC. The unperturbed frequencies are (c) $\Omega_{1}=5.384-i\,0.026$, (g) $\Omega_{2}=5.326-i\,0.027$ and (k) $\Omega_{3}=5.255-i\,0.004$.
There are many types of ARCs evident in Fig. 7. The TM-TE type ARC shown in Fig. 8 in the main text is one possible combination of mode polarizations. Crossings involving more than one mode also exist, one if which is shown in Fig. 16. It involves two TE and one TM polarized modes. Noticeably, only the upper and the lower branch exhibit mode hybridization at the ARC at $\delta=0.12$; this is also visible in the crossing of imaginary components of $\Omega$ in Fig. 16(b). The middle branch experiences an enhanced lifetime, which is caused by the mode pattern resembling a mode with higher $\eta$ than the unperturbed mode. Modes with high $\eta$ generally have longer lifetimes. The behavior of an ARC involving $n$ modes is described by an $n\times n$ Hamiltonian [40, 75].
@@ Appendix E Husimi functions at the EP
(a) $\Omega_{1}=4.831-i\,0.033$
(b) $\Omega_{2}=4.834-i\,0.033$
Figure 17: Top row: Electric (red) and magnetic (blue) fields. Bottom row: Associated average Husimi functions close to the EP at $\delta=0.222$} and n=2.435{n=2.435}.